(1096) Problem 7: A of radius r- e.38 m connects to the original by a thin conducting wite. The spheres are separated by an enormous dastance compared sphere 80% atan0 acotano sinb0 Partial Hint 0)-2.053 X Attempts Remai
Answer #1

charhe floe fill takes placed till the potential of the spheres equlises

k ( q1 - q2) / r1 = k q2 / r2

q1 / r1 = q2 / r2 + q2 / r1

q1 = r1 * q2 * ( r1 + r2) / r1r2

q2 = r2 q1 / ( r1 + r2)

q2 = 3.7 * 0.38 / ( 0.38+ 0.49)

q2 = 1.616 uC

==========